OpenStudy (anonymous):

4(3/5)*7(1/4)=(667/20) how i keep getting 6440/20

OpenStudy (ash2326):

tabby it's \[ 4 \frac {3}{5}* 7 \frac{1}{4}\] so it's equal to \[ (4+ \frac {3}{5})*(7+\frac{1}{4})\] now we get \[( \frac{23}{5})*(\frac{29}{4})\] so we get \[ \frac{667}{20}\]

OpenStudy (ash2326):

Got it??

OpenStudy (anonymous):

how did you get 23 and 29

OpenStudy (ash2326):

\(4\frac{3}{5}\) and \(7\frac{1}{4}\) are mixed fractions, Do you know mixed fractions?

OpenStudy (ash2326):

\[4\frac{3}{5}=4 + \frac{3}{5}= \frac{4*5}{5}+\frac{3}{5}=\frac{20}{5}+\frac{3}{5}=\frac{23}{5}\] Similarly \[7 \frac{1}{4}= \frac{29}{4}\]

OpenStudy (anonymous):

yes i got it know thank you so much

OpenStudy (ash2326):

glad to help Tabby:)